A consumer spends all his income (in equilibrium) to purchase 50 units of x and 20 units of y when the price of x and y are Rs. 2 and Rs.5 per unit per unit respectively


1
Expert's answer
2021-08-29T16:54:11-0400

Complete Question:

A consumer spends all his income (in equilibrium) to purchase 50 units of x and 20 units of y when the price of x and y are Rs. 2 and Rs.5 per unit per unit respectively


(i) Write the equation of his budget line.


Solution

"BL:=XP_X+YP_Y=M\\Where, P_X=Price of X\\P_Y=Price of Y\\And, M=Income\\X=50\\Y=20\\P_X=2\\P_Y=5\\M=XP_X+YP_Y\\M=50\times2+20\times 5\\M=100+100\\M=200\\BL:XP_X+YP_Y=M\\X\times2+Y\times5=200\\2X+5Y=200"


Therefore, the equation of the budget line is 2X+5Y=200.